OpenStudy (1dfanatic):

A line passes through (7,2) and (11,-14). Write the equation of the line in general form. A. 4x + y = 26 B. 4x + y = 30 C. 4x - y = 1 D. 4x - y = 26

OpenStudy (anonymous):

firstly, you want to work out the gradient can you do so?

OpenStudy (1dfanatic):

Whats a gradient?

OpenStudy (anonymous):

its the slope

OpenStudy (anonymous):

okay so try using this equation \[slope = \frac{ y - y1 }{ x - x1 }\]

OpenStudy (1dfanatic):

7-11?

OpenStudy (anonymous):

yes thats the bottom part of the fraction, what do you think will go on the top?

OpenStudy (1dfanatic):

2-14

OpenStudy (anonymous):

close, its actually 2 - (-14)

OpenStudy (anonymous):

which becomes 2 + 14

OpenStudy (anonymous):

but now we can put the bottom and top of our fraction together to simplify, what would you get?

OpenStudy (1dfanatic):

-1/4

OpenStudy (anonymous):

not exactly... \[\frac{ 2--14 }{ 7 - 11 }\]

OpenStudy (1dfanatic):

then -4

OpenStudy (anonymous):

yup thats right and we can now sub this into the equation \[y - y1 = m(x - x1)\] so try this and post up what you get

OpenStudy (1dfanatic):

idk where to substitute it in..

OpenStudy (1dfanatic):

like i dont get the whole letters with math thing..

OpenStudy (kayders1997):

A point that was given to you

OpenStudy (kayders1997):

For the x1 and y1

OpenStudy (anonymous):

yea so remember your slope = m (so thats the -4 you just worked out) and (x1, y1) you can choose to either be (7,2) or (11, -14) (because both these points lie on the line

OpenStudy (1dfanatic):

-4(7-11)...? I honestly don't know..

OpenStudy (kayders1997):

Plug the point into that equation for x1 and y1 so your 7 would be you x1 and 2 would be you y1 and out the slope in

OpenStudy (kayders1997):

And just x and y with nothing has nothing put in it, it just says x and y

OpenStudy (anonymous):

just remember that we don't need to substitute anything in for x and y, only for x1 and y1 so the two possible sub ins you can do are..\[y - (2) = (-4)(x - (7))\] or \[y - (-14) = (-4)(x - (11))\]

OpenStudy (anonymous):

either way when you simplify you should get the same answer, you can choose to sub in whichever point you want in the form as i have shown above (but you would like to pick the first option because that way you are dealing with smaller numbers

OpenStudy (kayders1997):

Do you know what to do now

OpenStudy (1dfanatic):

No and i just give up because im still not understanding thanks for yalls help though

OpenStudy (kayders1997):

No dont give up what part is confusing you?

OpenStudy (kayders1997):

You can do it! :)

OpenStudy (anonymous):

yeah which bit don't you understand? its okay we can help you :)

OpenStudy (1dfanatic):

everything...math is so confusing to me no matter how much help i get..

OpenStudy (kayders1997):

So you have the point (x1,y1) it doesnt mattee which you choose so you have the point (7,2) so you 7 is your x1 and 2 is your y1

OpenStudy (kayders1997):

So now you plug that into your equation y-y1=m(x-x1) y1 and x1 is your point and y and x just stay the same, than plug in your slope too

OpenStudy (kayders1997):

You can do it 1Dfanatic I know you can